An amount of 0.450 mole of an unknown hydrocarbon was subjected to combustion. The process produced 980.0 mL of CO2 at STP.

a. Find the volume of CO2 formed at STP from one mole of the unknown hydrocarbon.

b. Compute the number of carbon atoms present in one molecule of the unknown hydrocarbon.


1
Expert's answer
2022-01-12T09:20:02-0500

Molar mass of hydrocarbon=28.06

28.06×0.45=12.627

= 980/12.627=77.61ml